# Quantum information with quantum-like bits

### v1

@@ III Quantum-like gates
In this section, we discuss the fundamental principles of QL information processing. Specifically, we demonstrate how the emergent states of QL resources can be transformed consistently to any arbitrary quantum gate. We then outline a perspectives on how to translate these transformations onto operations on the edges of the graph resources. This is particularly relevant for the future computational and experimental implementation of QL algorithms.
@@
denotes the standard Hadamard (H) gate. ?? acts on the factorized states as
$U_{\mathrm{cb}}\psi_{\bm{\sigma}}$ $=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\begin{pmatrix}\frac{1}{2}[1+\sigma^{(q)}]1_{N_{G}}\\ \frac{1}{2}[1-\sigma^{(q)}]1_{N_{G}}\end{pmatrix},$ (III.3)
@@
The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$, where
$U_{\mathrm{cb}}\psi_{+}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{-}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix}.$ (III.5)
@@
$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (III.6)
@@
$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.7a)
$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.7b)
$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.7c)
$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.7d)
@@
$U_{\mathrm{H}}\psi_{\pm}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{\pm}=\frac{1}{\sqrt{2}}(\psi_{+}\pm\psi_{-}).$ (III.9)
@@
$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}},$ (III.10)
@@
$U_{\mathrm{CN}}\psi_{+\pm}=\psi_{+\pm},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{-\pm}=\psi_{-\mp}.$ (III.11)
Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis. We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??. In future work, we aim to design a bijective map $\psi\mapsto\mathcal{R}_{\psi}$ between QL states and resources. This approach enables the mapping of any gate transformation on states, $\psi\mapsto U_{\mathrm{g}}\psi=\psi_{\mathrm{g}}$, onto operations on cluster of edges of the corresponding resources, $\mathcal{R}_{\psi}\mapsto\mathcal{R}_{\psi_{\mathrm{g}}}$. This provides a clear numerical and experimental framework for QL computing.
In the next ?? we shift the discussion towards the benefits provided by a suitable measure theory designed to add “quantumness” to an otherwise classical network. This framework fulfills premises which can be key for a future quantification of QL advantage, as originally proposed in Ref. Khrennikov, 2019.
@@ IV Quantum-like measurements
In this section, we discuss criteria that allows to introduce a notion of QL interference of probabilities in our approach. This highlights that genuine quantum mechanical systems are not strictly required to build QL stochastic models provided the notions of states, observables and measurement are defined according to a given set of prescriptions.
@@ IV.1 Axiomatic definition
States can be expressed by linear combinations of elements of the factorized basis ??:
$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (IV.1)
@@
QL observables are represented by self-adjointed operators on $\mathcal{H}$. Their action on QL states, expressed by linear maps, defines a QL measurement.
Observables are expressed in term of linear combinations with structure
@@
and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$, as defined in ??. The QL measurement of $\mathcal{F}$ on a state $\psi$ is defined by the map $\psi\mapsto\mathcal{F}\psi$.
@@
the QL probability of collapsing onto the subspace spanned by the eigenvector $f_{\bm{\sigma}}$, with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$. ?? is expressed in terms of the projector
@@
the QL conditional probabilities associated to the consecutive measures of $\mathcal{F}$ and $\mathcal{G}$ on a state $\psi$.
@@ IV.2 Quantum probabilities in a QL model
As discussed in Ref. Khrennikov, 2021, an important feature of the QL probability formula ?? is that it does not require the calculation of exponentially time-consuming joint probability distributions. Instead, these are necessary for the classical updates described by ??. This observation is connected to the foundation of quantum advantage. We aim to explore this relevant property in future test of QL algorithms.
V Conclusions
@@
In this paper, we expanded upon the general approach derived in Refs. Scholes, 2023, 2024; Scholes and Amati, 2024 to define and manipulate QL states. In ?? we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs. We introduced the Hilbert space for an arbitrary number of QL bits, and we derived explicitly the Bell states in dimension two. We then proposed in ?? a notion of quantum gates applicable to QL states, and we provided concrete examples of their action. Finally, we discussed in ?? a list of measurement axioms which allow to observe QL interference effects in our approach.
@@
In the paper we introduced notions of entanglement, interference, and QL probabilities which avoid the expensive calculation of joint distributions. In future work we plan to investigate how each of these features contributes to achieving a “QL advantage” over strictly classical information processing. The ultimate test of the approach proposed in this work will be to implement benchmark algorithms, and evaluate their time efficiency and resource cost. Bharti et al. (2022)
In the last decades, advanced error correction techniques have been developed to minimize loss of information and decoherence in quantum processors. Cai et al. (2023); Campbell (2024) Similarly, we can ask how to prevent and minimize errors in our framework, e.g. in presence of impurities and irregularities in the structure of the graphs. In this regard, we note that a foundational property of synchronizing graphs is that their robustness against uncontrolled noise and dissipation increases with size. Scholes (2023); Scholes and Amati (2024); Anwar et al. (2023) Therefore, we conjecture that error correction is achievable in our approach by simply increasing the number of classical nodes of the graphs. We aim to further explore this concept in future work, to provide bounds of fidelity of gate transformations as a function of the size of the system and strength of external noise.
@@
From an experimental standpoint, different options can be explored to implement synchronizing resources for QL information processing. Insulator-to-metal-transition devices and spin-torque oscillators are viable candidates, as they allow for both distributed and parallel computing, while ensuring good robustness to noise and fault tolerance. Raychowdhury et al. (2019) In future work, we aim to further analyze platforms to experimentally implement QL resources, and the properties required by different approaches for optimal network implementation and manipulation.
@@
Finally, it can be a fascinating topic for future research to connect QL information processing introduced here to other information theories on human decision-making and psychology derived in the framework of contextual probability models. This idea has been extensively discussed in the literature of QL probabilistic models.Khrennikov (2016); Ozawa and Khrennikov (2020); Scholes (2024); Widdows, Rani, and Pothos (2023) Implementing our formalism on brain-like neuromorphic circuits Pal et al. (2024) can foster theoretical and experimental investigations in this framework.
VI Acknowledgments

### v4

@@ III Quantum-like gates
In this section, we discuss the fundamental principles of QL information processing. Specifically, we demonstrate how the emergent states of QL resources can be transformed consistently to any arbitrary quantum gate. We then study how to translate these transformations into operations on the edges of the graph resources. This is particularly relevant for the future computational and experimental implementation of QL algorithms.
@@
denotes the standard Hadamard (H) gate. The action of $U_{\mathrm{cb}}$ becomes clear for ${N_{\mathrm{QL}}}=1$. Here,
$U_{\mathrm{cb}}\psi_{\uparrow}=\begin{pmatrix}1_{N_{G}}\\ 0_{N_{G}}\end{pmatrix},\hskip 14.22636ptU_{\mathrm{cb}}\psi_{\downarrow}=\begin{pmatrix}0_{N_{G}}\\ 1_{N_{G}}\end{pmatrix},$ (III.3)
@@
$U_{\mathrm{g}}=U_{\mathrm{cb}}^{-1}(V_{\mathrm{g}}\otimes\mathbb{1}_{{N_{G}}})U_{\mathrm{cb}}.$ (III.5)
@@
$U_{\mathcal{I}}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.6a)
$U_{X}$ $=\begin{pmatrix}\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\\ \mathbb{0}_{N_{G}}&-\mathbb{1}_{N_{G}}\end{pmatrix},$ (III.6b)
$U_{\mathrm{Y}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&i\mathbb{1}_{N_{G}}\\ -i\mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.6c)
$U_{\mathrm{Z}}$ $=\begin{pmatrix}\mathbb{0}_{N_{G}}&\mathbb{1}_{N_{G}}\\ \mathbb{1}_{N_{G}}&\mathbb{0}_{N_{G}}\end{pmatrix},$ (III.6d)
@@
$U_{\mathrm{H}}\psi_{[\uparrow\downarrow]}=\left(V_{\mathrm{H}}\otimes\mathbb{1}_{{N_{G}}}\right)\psi_{[\uparrow\downarrow]}=\frac{1}{\sqrt{2}}(\psi_{\uparrow}\pm\psi_{\downarrow}).$ (III.8)
@@
$U_{\mathrm{CN}}$ $=\frac{1}{2}(U_{\mathcal{I}}+U_{\mathrm{Z}})\otimes U_{\mathcal{I}}+\frac{1}{2}(U_{\mathcal{I}}-U_{\mathrm{Z}})\otimes U_{\mathrm{X}},$ (III.9)
@@
$U_{\mathrm{CN}}\psi_{\uparrow[\uparrow,\downarrow]}=\psi_{\uparrow[\uparrow,\downarrow]},\hskip 28.45274ptU_{\mathrm{CN}}\psi_{\downarrow[\uparrow,\downarrow]}=\psi_{\downarrow[\downarrow,\uparrow]}.$ (III.10)
Note that, since a universal set of gates can be constructed by combining one-QL-bit gates and the CNOT, we can limit our analysis to this set of operations. Barenco et al. (1995)
Up to this point, we have focused exclusively on the action of QL gates on states $\psi_{\bm{\sigma}}$ of the factorized basis. We can infer from those the transformation of any arbitrary state $\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\gamma_{\bm{\sigma}}\psi_{\bm{\sigma}}$, $\gamma_{\bm{\sigma}}\in\mathbb{C}$, thanks to linearity of the map ??. From an operational standpoint, it is however preferable to translate these state operations onto transformations applicable directed to the graph resources. This feature is key to translate our gate formalism into practical algorithms for classical computing resources and for experimental platforms of synchronizing electrical circuits. Ochs, Michaelis, and Roggendorf (2019) In this regard, we observe that the network transformation
$\mathcal{R}\mapsto\mathcal{R}_{\mathrm{g}}=U_{\mathrm{g}}\mathcal{R}U_{\mathrm{g}}^{\dagger}$ (III.11)
transforms all eigenstates of a resource consistently to the gate $U_{\mathrm{g}}$ as defined in ??. This can be inferred from the spectral decomposition of ??, given by
$\mathcal{R}=\sum_{i=1}^{\tilde{N}}\lambda_{i}\psi_{i}\otimes\psi_{i}^{\dagger}.$ (III.12)
Here, $\tilde{N}=(2{N_{G}})^{N_{\mathrm{QL}}}$ denotes the total number of eigenstates of the resource $\mathcal{R}$, including both the emergent and the random states.
In the next ?? we shift the discussion towards the benefits provided by a suitable measure theory designed to add “quantumness” to classical synchronizing networks. This framework fulfills premises which can be key for a future quantification of QL advantage, as originally proposed in Ref. Khrennikov, 2019.
@@ IV Quantum-like measurements
In this section, we discuss criteria that allow to introduce a notion of QL interference of probabilities in our approach. This analysis highlights that genuine quantum mechanical systems are not strictly required to build QL stochastic models provided the notions of states, observables and measurement are defined according to a given set of prescriptions.
@@ IV.1 Axiomatic definition
States can be expressed by linear combinations of elements of the factorized basis in ??:
$\psi=\sum_{\bm{\sigma}\in\bm{\mathcal{S}}}\alpha_{\bm{\sigma}}\psi_{\bm{\sigma}}.$ (IV.1)
@@
QL observables are represented by self-adjointed operators on $\mathcal{H}$. Their action on QL states, expressed by a linear map, defines a QL measurement.
Observables are defined in terms of linear combinations with structure
@@
and fulfill the self-adjointness condition $\mathcal{F}^{\dagger}=\mathcal{F}$ [??]. The QL measurement of $\mathcal{F}$ on a state $\psi$ is defined by the map $\psi\mapsto\mathcal{F}\psi$.
@@
the QL probability of collapsing onto the subspace spanned by the eigenvector $f_{\bm{\sigma}}$ with eigenvalue $\lambda^{\mathcal{F}}_{\bm{\sigma}}$, after measuring $\mathcal{F}$ on $\psi$. ?? is expressed in terms of the projector
@@
the QL conditional probabilities associated to the consecutive measures of $\mathcal{F}$ and $\mathcal{G}$ on $\psi$.
@@ IV.2 Quantum probabilities in a QL model
As noted in Ref. Khrennikov, 2021, a significant feature of the QL probability formula ?? is that it bypasses the need for calculating joint probability distributions, which are typically exponential in computational complexity. In contrast, such calculations are required for the classical updates outlined in ??. This distinction relates directly to the foundational principles of quantum advantage. We intend to further investigate this property in future testing of QL algorithms.
V End-to-End quantum-like information processing and measurement
For genuine quantum resources, the primary source of error during information processing and measurement arises from the presence of uncontrolled noise. This can result from a variety of sources, including external electromagnetic fields, undesired interactions between qubits, fabrication defects, and imperfections in control protocols. Hou et al. (2016); Bengtsson et al. (2024); Funcke et al. (2022); Gupta and Nene (2021); Resch and Karpuzcu (2021) To account for these sources of error the final state of a quantum register after computation requires multiple measurements, which can lead to a significant overhead in the computing protocols. Likewise, error mitigation techniques and decoding protocols can significantly affect scalability and performance. Bravyi et al. (2021); Ali et al. (2024). In the following, we analyze both the accuracy and computational complexity of information processing and measurement within the QL formalism.
Given that QL resources are intrinsically classical, noise from external sources of dissipation does not appear to be a fundamental limitation in our approach. Instead, the sources of errors are intrinsic to the formalism itself, impacting on an equal footing information processing and final measurement. Here, we present an analysis of these errors by examining the overall workflow of QL information protocols from end to end.
Let us consider a generic QL algorithm consisting of a set of $M$ gates $\mathrm{g}=\{g_{1},\cdots,g_{M}\}$. This corresponds to the chain of transformations $U_{\bm{\mathrm{g}}}=\prod_{j=M}^{1}U_{g_{j}}$, as defined in ??. An exact approach to QL computing would involve applying this sequence of gates to transform via ?? the initial register $\mathcal{R}$ [??]. This consists of the Cartesian product of ${N_{\mathrm{QL}}}$ factors, as illustrated in panel a) of ??. Hence, storing and processing information following this exact strategy involves an exponential overhead with the number of edges in the total resource. A key insight we aim to leverage to reduce the complexity of this approach is as follows: if we can develop approximate QL protocols that maintain the structure of the Cartesian product, the exponential operations required for information processing and measurement could be circumvented entirely. We refer to this approach as the Cartesian-product approximation (CPA). This strategy enables to implement gate operations by transforming only a small subset of the edges of the total resources. In fact, within the CPA, information can be conveniently stored and processed with linear complexity in ${N_{\mathrm{QL}}}$ [see panel b) of ??].
Firstly, we assess the degree to which the CPA holds within our approach and quantify the associated errors relative to exact QL computing. Conveniently, the Cartesian-product structure is exactly preserved by all one-QL-bit gates , defined by
$W^{(p)}_{\mathrm{g}}=\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes U^{(p)}_{\mathrm{g}}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)},$ (V.1)
acting locally on a single resource $p\in\{1,\cdots,{N_{\mathrm{QL}}}\}$. ?? transforms the initial graph resource ?? as
$\displaylimits_{q=1}^{N_{\mathrm{QL}}}\mathcal{G}^{(q)}\;\begin{subarray}{c}U^{(p)}_{\mathrm{g}}\\ \longmapsto\end{subarray}\;\mathcal{G}^{(1)}\square\cdots\square\mathcal{G}^{(p)}_{\mathrm{g}}\square\cdots\square\mathcal{G}^{({N_{\mathrm{QL}}})},$ (V.2)
where $\mathcal{G}^{(p)}_{\mathrm{g}}$ is the gate transformation of the graph $\mathcal{G}^{(p)}$. To extend the approach to a universal set of gates, we analyze how effectively the CPA approximation holds for the CNOT operation $U_{\mathrm{CNOT}}^{(q,p)}$, applying the Pauli-X gate to the $p$-th QL bit depending on the state of the control $q$. As shown in ??, this gate conserves the Cartesian product only approximately, namely
$\displaylimits_{q=1}^{N_{\mathrm{QL}}}\mathcal{G}^{(q)}\;$ $\begin{subarray}{c}U_{\mathrm{CNOT}}^{(q,p)}\\ \longmapsto\end{subarray}\;\mathcal{G}^{(1)}\square\cdots\square\tilde{\mathcal{G}}^{(p)}\square\cdots\square\tilde{\mathcal{G}}^{(q)}\square\cdots\square\mathcal{G}^{({N_{\mathrm{QL}}})}$
$\quad+\mathcal{O}(l/k).$ (V.3)
?? introduces an error which depends on the ratio between the valencies $l$ and $k$ of the elementary subgraphs in ??, respectively $C$ and $A_{i}$. This result indicates that the optimal choice for the parameters defining the subgraphs are
$l=1,\hskip 14.22636ptk={N_{G}}-1,\hskip 14.22636pt{N_{G}}\gg 1.$ (V.4)
In future work, we aim to investigate the propagation of error in ?? as several gates are concatenated to construct QL algorithms.
An illustration of the final measurement process after information processing is shown in panel c) of ??. Let us firstly consider factorized measurement operators, with structure
$\mathcal{F}_{{\bm{g}}^{\prime}}=\bigotimes_{q=1}^{N_{\mathrm{QL}}}U^{(q)}_{g^{\prime(q)}}.$ (V.5)
Conveniently, even at the measurement state we can avoid to explicitly calculate the exponentially expensive Cartesian product for this common class of observable operators. In fact, as shown in panel c) of ??, local unitary transformations applied to one-QL-bit resources provide complete information on the measurement probabilities ??. For example, a local measurement of the population difference of the $q$-th QL bit acts on the final local emergent states as
$\psi^{(q)}_{\sigma^{(q)},\bm{\mathrm{g}}}$ $\mapsto U_{\mathrm{Z}}^{(q)}\psi^{(q)}_{\sigma^{(q)},\bm{\mathrm{g}}},$ (V.6a)
[…]
The more general case of non-factorized observables involves the sum over $M$ terms ?? expressed by
$\mathcal{F}_{\bm{g}_{1}^{\prime},\cdots,\bm{g}_{M}^{\prime}}=\sum_{m=1}^{M}\mathcal{F}_{\bm{g}^{\prime}_{m}}.$ (V.7)
The action of ?? breaks the structure of the Cartesian product, leading to a measurement cost that scales exponentially with the number of QL bits. It is worth noting that in a broad class of quantum algorithms, including Grover’s search Giri and Korepin (2017) and variational quantum eigensolvers, Tilly et al. (2022) the measurement process involves exclusively factorized operators, with structure as in ??. In these cases, this local measurement process can be applied with no limitations. We aim to compile a comprehensive list of the potential applications of our approach in future work.
Figure 5: a) Representation of the initial register, prior to information processing. The number of vertices of the total graph scales exponentially with ${N_{\mathrm{QL}}}$, the number of QL-bits. b) Information processing implemented within the CPA, as discussed in the main text. Here, we store and transform the single resources individually, with a favorable linear scaling with size. c) Illustration of the final measurement process. This involves the local diagonalization of the gate-transformed resources, to obtain ${N_{\mathrm{QL}}}$ sets of single-particle emergent states. Factorized measurements operators can be implemented locally on these resources, without calculating the tensor product associated to the final global emergent state.
VI Conclusions
@@
In this paper, we expanded upon the general approach derived in Refs. Scholes, 2023, 2024; Scholes and Amati, 2024 to define and manipulate QL states. In ?? we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs. We introduced the Hilbert space for an arbitrary number of QL bits, and we derived explicitly the Bell states in dimension two. In ?? we proposed a notion of quantum gates applicable to QL states, and we provided concrete examples of their action. We discussed in ?? a list of measurement axioms which allows to observe QL interference effects in our approach. Finally, we outlined in ?? an efficient end-to-end workflow for building QL algorithms based on a controlled approximation for two-QL-bit gates.
@@
In the paper we introduced notions of entanglement, interference, and QL probabilities which avoid the expensive calculation of joint probability distributions in the measurement process. In future work we plan to investigate how each of these features contributes to achieving a “QL advantage” over strictly classical information processing. The ultimate test of the approach proposed in this work will be to implement the pipeline proposed in ?? (and related ??) to study QL algorithms, and evaluate their time efficiency and resource demand. Bharti et al. (2022)
@@
In this paper, we introduced a universal set of gates comprising single QL-bit terms and the CNOT operation. Our controlled approximation is designed to preserve the Cartesian product structure by selectively neglecting a small contribution within the CNOT gate that would otherwise lead to symmetry breaking. Preserving this contribution would result in an exponential increase in complexity, posing significant challenges for both computational and experimental resources. An open question remains as to whether directly implementing an approximation of many-QL-bit gates, rather than decomposing them into one- and two-body elementary terms, could yield even greater accuracy. While any local decomposition is exact in standard unitary quantum computing, this is in general not the case within our QL approach. Additionally, alternative choices of universal sets may provide greater computational advantages. In future work, we aim to explore the impact of these different scenarios within our framework.
From an experimental standpoint, various possibilities can be considered for implementing synchronizing resources in QL information processing. The research on oscillator-based computing, as reviewed in Ref. Csaba and Porod, 2020, offers a solid foundation for this purpose. Within this framework, insulator-to-metal-transition devices and spin-torque oscillators emerge as promising candidates, supporting distributed and parallel computing while ensuring resilience to noise and fault tolerance.Raychowdhury et al. (2019) We plan to further investigate platforms for the experimental implementation of QL resources, focusing on the properties needed by various approaches to achieve optimal network design and control.
In recent years, memcomputing has emerged as an innovative computational framework that exploits the intrinsic time-nonlocal properties of classical dynamical systems to enable a terminal-agnostic framework to implement non-Boolean logic. Di Ventra and Traversa (2018) This approach, which allows the computation to occur in memory rather than through sequential logic operations, could be exploited by recasting classical synchronizing dynamics leading to emergent states into time-nonlocal equations of motion. Amati, Meyer, and Schilling (2019) The classical Nakajima–Mori–Zwanzig formalism, which captures the influence of past states on open classical systems, offers a strategy to incorporate memory effects directly into the QL formalism, potentially enhancing stability and coherence.
@@
Finally, it can be a fascinating topic for future research to connect QL information processing introduced here to other information theories on human decision-making and psychology derived in the framework of QL contextual probability models. Khrennikov (2016); Ozawa and Khrennikov (2020); Scholes (2024); Widdows, Rani, and Pothos (2023) Implementing our formalism on brain-like neuromorphic circuits Pal et al. (2024) can support theoretical and experimental investigations in this framework.
VII Acknowledgments